25.      Which of the following is management’s challenge when setting transfer prices?

a.

Ensuring the buyer has goal congruence with respect to the organization’s goals.

b.

Ensuring the seller has goal congruence with respect to the organization’s goals.

c.

Ensuring either the buyer or the seller, but not both, has goal congruence with respect to the organization’s goals.

d.

Ensuring both the buyer and seller have goal congruence with respect to the organization’s goals.

26.      Which of the following transfer pricing procedures maximizes the company’s profit by transferring at the differential outlay cost to the selling division plus the opportunity cost to the company of making the internal transfers?

a.

effective transfer pricing rule.

b.

economic transfer pricing rule.

c.

efficient transfer pricing rule.

d.

None of the answers is correct.

27.      What is generally considered the best transfer pricing basis when there is a competitive market for the product and market prices are readily available?

a.

Market price-based transfer pricing

b.

Variable cost-based transfer pricing

c.

Fixed price-based transfer pricing

d.

Fixed cost-based transfer pricing

Colonial Computing Systems

Colonial Computing Systems manufactures and sells various computer products and has two decentralized divisions: (1) Production and (2) Marketing. The Marketing Division has always purchased a particular mouse from Production at $55 per unit. The Production Division is considering raising the price to $60 per unit. The Production Division's costs related to the mouse production is as follows:

Variable costs per unit:

$55

Monthly fixed costs:

$12,000

The Marketing Division handles the promotion and distribution of the mouse purchases from the Production Division and sells each mouse for $100. Marketing Division incurs monthly fixed costs of $5,000. Marketing Division sells 1,500 units per month. Marketing Division can buy the same mouse from outside suppliers for $60. If the Marketing Division purchases the mouse from outside suppliers, the facilities the Production Division uses to manufacture the mouse would remain idle.

28.      Refer to Colonial Computing Systems. The Production Division is operating below capacity because of weak global demand for the product. What should be the mouse transfer price between the Production Division and Marketing Division in order for Colonial to optimize profits?

a.

$  60

b.

$  55

c.

$  50

d.

$100
Answer #1

25. d. while setting Transfer Prices ; Management has to ensure that both buyer and seller have goal congruence with respect to the organisation's goals.

26. b. GENERAL ECONOMIC TRANSFER PRICING RULE maximizes the company’s profit by transferring at the differential outlay cost to the selling division plus the opportunity cost to the company of making the internal transfers.

27. a. when there is a competitive market for the product and market prices are readily available then MARKET PRICE-BASED TRANSFER PRICING is generally considered the best transfer pricing basis.

28. b. 55$

Calculation:

Colonial Computing Systems:

- If Production division transfers to marketing division at 55:

Prod Division:

Sales 1500@55= 82500

Less: Variable Cost 1500@55=(82500)

Less: Fixed Cost=(12000)

Total= (12000)

Marketing Division:

Sales 1500@100=150000

Less: Internal Purchase (82500)

Less: Fixed Cost (5000)

Total= 62500 ;

overall to the colonial = (12000)+62500=50500

- If Production division transfers to marketing division at 60:

Prod Division:

Sales 1500@60= 90000

Less: Variable Cost 1500@55=(82500)

Less: Fixed Cost=(12000)

Total= (4500)

Marketing Division:

Sales 1500@100=150000

Less: Internal Purchase (90000)

Less: Fixed Cost (5000)

Total= 55000 ;

overall to the colonial = (4500)+55000=50500

From the above calculation we can say that Transfer price doesn't affect the overall profit or loss of the enterprise (colonial). same in both cases "50500".

- If marketing Division Purchase mouses from outside suppliers @ 60

Prod Division:

Fixed Cost= (12000)

Marketing Division:

Sales 1500@100=150000

Less: Outside Purchase 1500@60=(90000)

Less: Fixed Cost (5000)

Total= 55000

overall to the colonial = (12000)+55000=43000

So from the above calculation we are having income of only 43000 if Marketing division purchase from outside so it should purchase from production division only.

Calculation of Transfer price in order to Maximize Colonial's profit:

If transfer @ 55-

Sales 1500@100=150000

Less: variable Cost 1500@55=82500

Contribution=67500

If transfer @ 60-

Sales 1500@100=150000

Less: variable Cost 1500@60=90000

Contribution=60000

CONCLUSION:- Contribution to Colonial will be more if Transfer Price is 55 so Transfer Price should be 55/-.